# Improving Reinforcement Learning from Human Feedback Using Contrastive Rewards

### 12 Mar 2024 v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is the mainstream paradigm used to align large language models (LLMs) with human preferences. Yet existing RLHF heavily relies on accurate and informative reward models, which are vulnerable and sensitive to noise from various sources, e.g. human labeling errors, making the pipeline fragile. In this work, we improve the effectiveness of the reward model by introducing a penalty term on the reward, named as contrastive rewards. Our approach involves two steps: (1) an offline sampling step to obtain responses to prompts that serve as baseline calculation and (2) a contrastive reward calculated using the baseline responses and used in the Proximal Policy Optimization (PPO) step. We show that contrastive rewards enable the LLM to penalize reward uncertainty, improve robustness, encourage improvement over baselines, calibrate according to task difficulty, and reduce variance in PPO. We show empirically contrastive rewards can improve RLHF substantially, evaluated by both GPTs and humans, and our method consistently outperforms strong baselines.
@@ 4.1 Setup
[…]
@@
  • Human-assisted Evaluation: We also engage the support of three individuals to annotate samples in cases where GPT-4 yields inconsistent judgments or declares a tie. However, we adopt GPT-4’s judgment if it consistently deems one answer superior to the other. Specifically, for each sample, we gather three annotations, and the final evaluation is determined by the majority vote among these annotations. To ensure the quality of human annotation, 30% of the labeled samples are conducted random examinations during each verification period. We only incorporate annotations when the annotator’s accuracy on our gold standard exceeds 90% during each verification period. If the accuracy falls below this threshold, the annotations are re-sampled until the requirement is met. The annotation rules and prompts used for GPT-4 evaluation can be found in the Appendix D.
@@ 4.2 Implementation Details
[…]
  • Reward Model Training: We train the reward model for all datasets initialized from the SFT model. We train the reward models for up to three epochs and select the model that achieves the minimum loss on the validation dataset.
  • RL optimization: We use prompts from the training dataset for training and partition the prompts in the validation dataset into two segments – one for validation and the other for testing. We select the best model based on the highest reward attained on the validation dataset.
@@
We use $\lambda r_{x,y}^{\rm RL}$ as the final reward for policy optimization.
Table 2: Win rate evaluated by third-party RM, UltraRM.
Table 3: Win rate evaluated by the third-party RM, PairRM.
@@
Table 3: Win rate evaluated by the third-party RM, PairRM.
@@ 4.4 Ablation Studies
We perform a series of ablations studies to better understand our method.
@@
Figure 2: Distribution of reward offsets $\Delta r$ ($\Delta r=r_{{x,y_{\text{highs}}}}-r_{{x,y_{\text{lows}}}}$). Distributions with the legend “lows” and “highs” represent the low-reward group and the high-reward group respectively.
@@
Table 5 presents the evaluation results on MT-Bench, capturing the average performance of the chatbot’s capabilities across 8 different dimensions. Leveraging contrastive rewards, i.e., Mistral-7B-CR, consistently outperforms the baseline models. We also include results from several open-source models alongside our methods for comparison. Notably, on MT-Bench, the model fine-tuned by RLHF-CR has surpassed the performance of Llama-70B-chat with a big margin (6.86 MT Score). For models other than Mistral, we directly utilize the MT score from the public leaderboard, excluding the 1st and 2nd results in Table 5. Detailed results in different dimensions are presented in Appendix B.

### 14 Mar 2024 v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) is the mainstream paradigm to align large language models (LLMs) with human preferences. Yet existing RLHF heavily relies on accurate and informative reward models, which are vulnerable and sensitive to noise from various sources, e.g. human labeling errors, making the pipeline fragile. In this work, we improve the effectiveness of the reward model by introducing a penalty term on the reward, named contrastive rewards. Our approach involves two steps: (1) an offline sampling step to obtain responses to prompts that serve as baseline calculation and (2) a contrastive reward calculated using the baseline responses in the Proximal Policy Optimization (PPO). We show that our contrastive rewards enable the LLM to penalize reward uncertainty, improve robustness, encourage improvement over baselines, calibrate according to task difficulty, and reduce variance in PPO. We also empirically demonstrate contrastive reward can improve RLHF substantially, evaluated by both GPTs and humans, and it consistently outperforms strong baselines.
@@ 4.1 Setup
  • Third-party Reward Model: In line with prior research (Eisenstein et al., 2023; Coste et al., 2023), we utilize public third-party reward models as evaluators. Specifically, we employ the well-established UltraRM-13B (Cui et al., 2023) and PairRM (Jiang et al., 2023) for evaluation. Both reward models are trained on the UltraFeedback dataset [^1] https://huggingface.co/datasets/openbmb/UltraFeedback, a large-scale, high-quality, and diversified preference dataset that has demonstrated effectiveness by various robust open-source models (Tunstall et al., 2023; Cui et al., 2023). More importantly, the majority of all three datasets we utilized are included in UltraFeedback, featuring refined high-quality annotations. Consequently, they are capable of providing accurate and convincing evaluation results. To compare the two models, we utilize the third-party reward models to score the responses generated by the two models in the test dataset, considering the model with the higher score as the winner. We then report both the average reward and win rate as determined by these two robust third-party reward models.
@@
  • Human-assisted Evaluation: We also engage the support of three individuals to annotate samples in cases where GPT-4 yields inconsistent judgments or declares a tie. We only adopt GPT-4’s judgment if it consistently deems one answer superior to the other. Specifically, for each sample, we gather three annotations, and the final evaluation is determined by the majority vote among these annotations. To ensure the quality of human annotation, 30% of the labeled samples are conducted random examinations during each verification period. We only incorporate annotations when the annotator’s accuracy on our gold standard exceeds 90% during each verification period. If the accuracy falls below this threshold, the annotations are re-sampled until the requirement is met. The annotation rules and prompts used for GPT-4 evaluation can be found in the Appendix D.
@@ 4.2 Implementation Details
  • Supervised Fine-tuning. All reward models and policy models undergo fine-tuning starting from Llama 7B (Touvron et al., 2023a) on the Supervised Fine-tuning (SFT) data across all datasets. This process aims at improving instruction-following capabilities for the task. For the dialogue task, i.e., Anthropic/HH-RLHF dataset and PKU dataset, they do not contain SFT data. Following previous work (Chiang et al., 2023), we utilize the ShareGPT dataset [^2] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, consisting of real human-interacted examples collected from ShareGPT.com, containing 821 million tokens for instruction fine-tuning. For the OpenAI/Summary task, which includes SFT data, we conduct supervised fine-tuning using this dataset.
  • Reward Model Training. We train the reward model for all datasets initialized from the SFT model. We train the reward models for up to three epochs and select the model that achieves the minimum loss on the validation dataset.
  • RL Optimization. We use prompts from the training dataset for training and partition the prompts in the validation dataset into two segments – one for validation and the other for testing. We select the best model based on the highest reward attained on the validation dataset.
@@
We use $\lambda\cdot r_{x,y}^{\rm RL}$ as the final reward for policy optimization.
Table 2: Win rate evaluated by third-party RM: UltraRM.
Table 3: Win rate evaluated by the third-party RM: PairRM.
@@
Table 3: Win rate evaluated by the third-party RM: PairRM.
@@ 4.4 Ablation Studies
We perform a series of ablations studies.
@@
Figure 2: Distribution of reward offsets $\Delta r=r_{{x,y_{\text{highs}}}}-r_{{x,y_{\text{lows}}}}$. Distributions with the legend “lows” and “highs” represent the low-reward group and the high-reward group respectively.
@@
Table 5 presents the evaluation results on MT-Bench, capturing the average performance of the chatbot’s capabilities across 8 different dimensions. Leveraging contrastive rewards, i.e., Mistral-7B-CR, consistently outperforms the baseline models. We also include results from several open-source models alongside our methods for comparison. Notably, on MT-Bench, the model fine-tuned by RLHF-CR has surpassed the performance of Llama-70B-chat with a big margin (6.86 MT Score). For models other than Mistral, we directly copy the MT score from the public leaderboard, therefore excluding the 1st and 2nd results in Table 5. Detailed results in different dimensions are presented in Appendix B.